# Structure-Enhanced Protein Instruction Tuning: Towards General-Purpose Protein Understanding

### v2

@@ Abstract
Proteins, as essential biomolecules, play a central role in biological processes, including metabolic reactions and DNA replication. Accurate prediction of their properties and functions is crucial in biological applications. Recent development of protein language models (pLMs) with supervised fine tuning provides a promising solution to this problem. However, the fine-tuned model is tailored for particular downstream prediction task, and achieving general-purpose protein understanding remains a challenge. In this paper, we introduce Structure-Enhanced Protein Instruction Tuning (SEPIT) framework to bridge this gap. Our approach integrates a noval structure-aware module into pLMs to inform them with structural knowledge, and then connects these enhanced pLMs to large language models (LLMs) to generate understanding of proteins. In this framework, we propose a novel two-stage instruction tuning pipeline that first establishes a basic understanding of proteins through caption-based instructions and then refines this understanding using a mixture of experts (MoEs) to learn more complex properties and functional information with the same amount of activated parameters. Moreover, we construct the largest and most comprehensive protein instruction dataset to date, which allows us to train and evaluate the general-purpose protein understanding model. Extensive experimental results on open-ended generation and closed-set answer tasks demonstrate the superior performance of SEPIT over both closed-source general LLMs and open-source LLMs trained with protein knowledge.
1 Introduction
@@
Proteins are large biomolecules and macromolecules, composed of one or more long chains of amino acid residues, playing pivotal roles in catalyzing metabolic reactions, DNA replication, and other significant biological processes (Anfinsen & Haber, 1961; Hartley, 1951). Generally, proteins are represented in two types of forms: one-dimensional (1D) sequence detailing the order of amino acids, and three-dimensional (3D) structure illustrating the spatial configuration of the protein. The 1D sequence of protein is generated by transcribing and translating a gene’s DNA sequence and folds into a specific 3D structure, and the 3D shape determines the protein’s properties and functions (Pei et al., 2024b). Conventional machine learning methods (Radivojac et al., 2013; Whisstock & Lesk, 2003) have achieved notable accuracy in protein property and function prediction via supervised learning. However, these methods are all task-specific as each model is restricted to predicting a particular property. The increasing need for comprehensive protein analysis in various fields, such as pathology and drug discovery (Guo et al., 2023), calls for the development of general-purpose protein understanding models capable of accurately predicting various protein properties and functions.
@@
In the fast-evolving era of large language models (LLMs), large efforts have attempted to leverage their capabilities of semantic understanding and complex reasoning to achieve general-purpose protein property and function prediction. Initially, some methods treat the 1D protein sequence as natural language input to LLMs (Wang et al., 2023b; Luo et al., 2023b; Taylor et al., 2022; Fang et al., 2024; Pei et al., 2024a). However, they focus on the learning of associations between protein sequences and their properties or functions using a small part of real-world protein sequences, which hinders the LLMs from generating reliable results for proteins at an evolutionary scale. To cope with this issue, ProtST (Xu et al., 2023) and ProteinCLAP (Liu et al., 2023) utilize pLMs pre-trained on evolutionary-scale protein databases as protein sequences encoders and leverage contrastive learning (Radford et al., 2021b) to train on protein-text paired data, thereby incorporating functional information from text with high-quality protein representations from pLMs. Unfortunately, these methods can be solely applied to prediction and retrieval tasks about proteins whereas real-world proteins with complex and diverse properties and functions require full understanding in an open-ended generation fashion instead of these specific tasks. Therefore, Prot2Text (Abdine et al., 2024), ProteinChat (Guo et al., 2023) and ProtT3 (Liu et al., 2024b) propose the protein-to-language generation via integrating protein sequence or structure encoding from pre-trained models.
@@
However, existing methods still show limitations in providing reliable general-purpose protein understanding for their application in scientific research. Firstly, although some studies have considered the determinant role of protein 3D structure on its properties and functions and used it as input, the proteins with directly usable 3D structural information is very rare. This leads to a situation where we need to learn the relationship between 1D sequences and functional information while relying on limited 3D information, making it challenging to provide reliable property and function predictions. Secondly, existing protein-related instruction datasets neglect 3D structures and have limited coverage of property and function types, which impedes the evaluation of model reliability and generalizability in general-purpose protein understanding tasks. Thirdly, the diversity of protein properties and functions poses a challenge for their accurate prediction. Using a single general-purpose model to predict a wide range and complex set of properties and functions is more challenging than fine-tuning multiple specialized models for different specific tasks.
@@
To address these challenges, we propose a generalized instruction tuning framework called Structure-Enhanced Protein Instruction Tuning (SEPIT) for general-purpose protein understanding. To comprehensively evaluate the reliability and generalizability of our models, we construct the largest protein instruction dataset to date which covers the most types of protein properties and functions, based on large-scale protein knowledge bases (Bairoch & Apweiler, 1997; Varadi et al., 2022). Before instruction tuning, in order to obtain a protein sequence/structure fused encoder that supports different types of protein input (1D or 1D&3D), we specially design a structure-aware module into pLMs. Then, we warm it up through protein-text contrastive learning and structure denoising, which provides a foundation to leverage a small amount of structural information for enhancing the understanding of large-scale sequence-only proteins. After that, based on the protein sequence/structure fused encoder, we design a two-stage protein instruction tuning pipeline to enable LLMs for general-purpose protein understanding. In stage 1, we instill basic understanding of proteins into the model through protein caption instructions. In stage 2, we initialize mixture of experts through upcycling (Komatsuzaki et al., 2023), which allows the model to learn more complex and diverse functions and properties, based on the basic understanding from stage 1, without increasing additional activated parameters. In summary, our contributions include: 1) By designing a structure-aware module and integrating it into pLMs, we enable the models to handle different types of protein inputs, thereby improving the quality of embedding over the vanilla sequence-only pLMs. 2) We construct the largest and most comprehensive protein instruction dataset to date, which addresses the lack of comprehensive dataset for general-purpose protein understanding. 3) We design a two-stage protein instruction tuning pipeline that enables a single model to learn a wide range and complex set of protein properties and functions. This is achieved by leveraging Mixtures of Experts (MoEs) built upon foundational knowledge. 4) Based on the proposed SEPIT framework and protein instruction dataset, we demonstrate the feasibility of enabling LLMs with the capability of general-purpose protein understanding.
2 Related Work
@@
In this section, we will present the related work pertinent to our study, focusing primarily on protein language models, multimodal instruction tuning and learning with 3D structural information.
Protein Language Models.
@@
Employing context-aware language models (Rosenfeld, 2000), protein sequences can be likened to sentences wherein amino acids serve as the elemental words. Through pre-training on databases containing hundreds of millions of such protein sequences (e.g., UniRef (Suzek et al., 2015), BFD (Steinegger et al., 2019; Steinegger & Söding, 2018)), pLMs enable effective modeling and prediction of protein structures and functions (Hu et al., 2022). In earlier works (Alley et al., 2019; Strodthoff et al., 2020; Heinzinger et al., 2019), LSTM and its variants (Hochreiter & Schmidhuber, 1997; Yu et al., 2019; Huang et al., 2015; Krause et al., 2016) were utilized to model the dependencies between residues in single protein sequences. With the rise of the Transformers architecture (Vaswani et al., 2017), Transformers-based pLMs emerged. ESM-1b (Rives et al., 2021), leveraging the Transformers architecture along with a masking strategy for pretraining, significantly enhances the prediction accuracy for mutational effects, secondary structure, and long-range contacts. After this, ProtTrans (Elnaggar et al., 2022) released two auto-regressive models (Dai et al., 2019; Yang et al., 2019) and four auto-encoder models (Devlin et al., 2018; Lan et al., 2019; Clark et al., 2020; Raffel et al., 2020) pre-trained on protein sequence databases. Beyond merely focusing on single protein sequences, MSA-Transformer (Rao et al., 2021) integrate multiple sequence alignments (MSA) of homologous proteins, provided a solid foundation for the success of AlphaFold2 (Jumper et al., 2021). Moreover, ESM-2 (Lin et al., 2023) further scaled up pLMs, achieving protein structure prediction performance comparable to previous works without utilizing MSA information, and significantly reduced inference overhead (Lin et al., 2022). Additionally, there were other studies that attempted to incorporate additional knowledge into the pre-training of protein sequences. For instance, ProteinBERT (Brandes et al., 2022) and OntoProtein (Zhang et al., 2022) integrated gene ontology (GO) information into the representations of protein sequences, enhancing the model’s understanding of protein functions. Although these pLMs can provide high-quality representations of proteins, they cannot generate natural language predictions about protein properties and functions.
Multimodal Instruction Tuning.
With the emergence of MLLMs such as GPT4 (OpenAI et al., 2024) and Genimi (Team et al., 2023), MLLMs had become a focal point of research. Initially, works like CLIP (Radford et al., 2021b), ALBEF (Radford et al., 2021a), VLMo (Radford et al., 2021c), SimVLM (Wang et al., 2021) emphasized exploring the cross-modal alignment between vision and language. Subsequently, based on modal alignment, Flamingo (Alayrac et al., 2022) and BLIP2 (Li et al., 2023) established bridges between visual encoders and LLMs using the Perceiver Resampler and the Q-Former, respectively. Following this, PaLM-E (Driess et al., 2023) introduced "multimodal sentences" as input, injecting real-world continuous sensor data into the LLMs in the form of language tokens, thereby endowing the model with a general multi-task capability. Additionally, efforts such as InstructBLIP (Dai et al., 2024), LLaVA (Liu et al., 2024a), MiniGPT4 (Zhu et al., 2023), mPLUGOwl (Ye et al., 2023), Qwen-VL (Bai et al., 2023), CogVLM (Wang et al., 2023a) applied the crucial instruction tuning technique from LLMs to MLLMs, enhancing the MLLMs’ ability to follow multimodal instructions. At the same time, they introduced innovations from the perspectives such as the construction of instruction datasets, training paradigms, and model design, which in turn refreshed the performance of MLLMs in a variety of visual-language downstream tasks (Yin et al., 2023). In this paper, we attempt to apply this paradigm to the protein domain, investigating the potential of endowing LLMs with general-purpose protein understanding capabilities.
Learning with 3D Structural Information.
Although pLMs pre-trained on protein sequences have been proven to be effective in numerous tasks, the protein structure is inherently a determinant of protein function (Anfinsen & Haber, 1961; Hartley, 1951). More effectively utilizing 3D structural information can help to understand proteins more comprehensively. To capture the impact of geometric positioning and interaction relationships among residues, a class of methods encoded 3D geometric information into rotation-invariant scalars, which was then processed through graph neural networks (GNNs) for message passing (Wu et al., 2022; Zhang et al., 2023a). For example, IEConv(Hermosilla et al., 2021) utilized a multi-graph to depict primary and secondary structures through covalent and hydrogen bonds and represented the tertiary structure with the spatial 3D coordinates of atoms. By blending intrinsic and extrinsic node distances and employing hierarchical pooling, it effectively perceived all three structural levels of proteins. GearNet went further by incorporating three types of directed edges (sequential edges, radius edges, and k-NN edges) into the graph, capturing information at various structural levels. On this basis, CDConv (Fan et al., 2023) utilized MLP to parameterize the kernel matrices, as opposed to employing distinct kernel matrices for varying edge types, which enabled a more flexible and efficient modeling of complex interactions between residues. Additionally, another class of methods sought to prevent the loss of 3D structural information by incorporating 3D rigid transformations into the network operations. This led to the development of geometric GNNs/Transformers characterized by SE(3) invariance and equivariance (Liao & Smidt, 2022; Liao et al., 2024; Fuchs et al., 2020; Zhou et al., 2023; Luo et al., 2023a; Wang et al., 2024). Representative examples of this approach, such as GVP (Jing et al., 2021) and EvoFormer (Jumper et al., 2021), were utilized by ESM-IF (Hsu et al., 2022), AlphaFold2 (Jumper et al., 2021), respectively. Furthermore, to leverage information from both evolutionary-scale protein sequences and the relatively limited protein structures, some other methods (Wang et al., 2022; Zhang et al., 2023a) attempted to build a bridge between these two domains.
3 Construction of Protein Instruction Dataset
@@
To endow LLMs with general-purpose protein understanding capabilities and evaluate their reliability and generalizability, in this paper, we construct a protein instruction dataset contains open-ended generation and closed-set answer tasks. For the open-ended generation subset, we mainly constructed it based on Swiss-Prot (Bairoch & Apweiler, 1997). We include almost all protein properties and functions contained therein (Function, Similarity, Subcellular location, Induction, Molecular Function, Biological Process, Cellular Component, Developmental Stage, Short Sequence Motif, Tissue Specificity, Activity Regulation, Pathway), and used ChatGPT (OpenAI et al., 2024) to aid in designing question templates based on the structured annotations. For the closed-set answer subset, we constructed it mainly based on the RCSB PDB (RCSB, 2024). We follow the data organized by previous researchers (Guo et al., 2023) and select parts of their proposed Q&A samples that are highly related to protein properties and functions, filtering out other samples related to metadata (e.g. discovery time and discovery methods). We have also sampled some examples related to Enzyme Commission (EC) and Gene Ontology (GO) predictions (Gligorijević et al., 2021) for inclusion in the closed-set answer subset. More detailed information about the dataset, including statistical information and examples, are shown in Appendix A.
@@
Compared to previously proposed protein-text related datasets (Xu et al., 2023; Fang et al., 2024; Wang et al., 2023b; Guo et al., 2023), our advantages are as follows: First, our dataset contains the most comprehensive set of instructions, covering almost all critical protein properties and function types found in databases (Bairoch & Apweiler, 1997). Second, our dataset includes the largest volume of instructions, comprising over 10 million instructions (with an additional 5 million supplementary instructions from TrEMBL). Third, our dataset incorporates structural information, offering the potential to enhance prediction reliability by leveraging this structural data. In summary, our dataset has the potential to support future research in the field of general-purpose protein understanding.
4 Structure-Enhanced Protein Instruction Tuning

### v3

@@ Abstract.
Proteins, as essential biomolecules, play a central role in biological processes, including metabolic reactions and DNA replication. Accurate prediction of their properties and functions is crucial in biological applications. Recent development of protein language models (pLMs) with supervised fine tuning provides a promising solution to this problem. However, the fine-tuned model is tailored for particular downstream prediction task, and achieving general-purpose protein understanding remains a challenge. In this paper, we introduce Structure-Enhanced Protein Instruction Tuning (SEPIT) framework to bridge this gap. Our approach incorporates a novel structure-aware module into pLMs to enrich their structural knowledge, and subsequently integrates these enhanced pLMs with large language models (LLMs) to advance protein understanding. In this framework, we propose a novel instruction tuning pipeline. First, we warm up the enhanced pLMs using contrastive learning and structure denoising. Then, caption-based instructions are used to establish a basic understanding of proteins. Finally, we refine this understanding by employing a mixture of experts (MoEs) to capture more complex properties and functional information with the same number of activated parameters. Moreover, we construct the largest and most comprehensive protein instruction dataset to date, which allows us to train and evaluate the general-purpose protein understanding model. Extensive experiments on both open-ended generation and closed-set answer tasks demonstrate the superior performance of SEPIT over both closed-source general LLMs and open-source LLMs trained with protein knowledge.
Large Language Models, Insturction Tuning, Protein
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadaisbn: 979-8-4007-1454-2/25/08doi: 10.1145/3711896.3737138ccs: Computing methodologies Natural language processingccs: Applied computing Bioinformatics
KDD Availability Link:
The code, datasets and model checkpoints of this paper has been made publicly available at https://doi.org/10.5281/zenodo.15510090.
1. Introduction
@@
Proteins are large biomolecules and macromolecules composed of one or more long chains of amino acid residues. They play pivotal roles in catalyzing metabolic reactions, facilitating DNA replication, and driving other essential biological processes (Anfinsen and Haber, 1961; Hartley, 1951). Generally, proteins are represented in two types of forms: a one-dimensional (1D) sequence, which specifies the order of amino acids, and a three-dimensional (3D) structure which depicts the spatial configuration of the protein. A protein’s 1D sequence is generated by transcribing and translating a gene’s DNA sequence and subsequently folds into a specific 3D structure. This 3D conformation determines the protein’s properties and functions (Pei et al., 2024b). Conventional machine learning methods (Radivojac et al., 2013; Whisstock and Lesk, 2003) have achieved notable accuracy in protein property and function prediction via supervised learning. However, these methods are all task-specific as each model is restricted to predicting a particular property. The growing demand for comprehensive protein analysis in fields such as pathology and drug discovery (Guo et al., 2023) highlights the need for general-purpose protein understanding models capable of accurately predicting a wide range of protein properties and functions.
@@
In the fast-evolving era of large language models (LLMs) (Zhao et al., 2025; Rosenfeld, 2000; Wu et al., 2025), significant efforts have been made to leverage their semantic understanding and complex reasoning capabilities to achieve general-purpose protein property and function prediction. Initially, some approaches treat the 1D protein sequence as natural language input to LLMs (Wang et al., 2024c; Luo et al., 2023b; Taylor et al., 2022; Fang et al., 2024; Pei et al., 2024a). However, they focus on learning associations between protein sequences and their properties or functions using only a limited subset of real-world protein sequences, which hinders the LLMs from generating reliable predictions at an evolutionary scale. To address this limitation, ProtST (Xu et al., 2023) and ProteinCLAP (Liu et al., 2025) utilize pLMs pre-trained on evolutionary-scale protein databases as protein sequence encoders and further employ contrastive learning (Radford et al., 2021a) on protein-text paired data to incorporate functional information from text with high-quality protein representations from pLMs. Unfortunately, these methods remain restricted to prediction and retrieval tasks. However, real-world proteins exhibit complex and diverse properties and functions that require a more comprehensive understanding in an open-ended generative manner rather than being confined to specific tasks. To bridge this gap, Prot2Text (Abdine et al., 2024), ProteinChat (Guo et al., 2023) and ProtT3 (Liu et al., 2024b) propose protein-to-language generation by integrating protein sequence or structure encoding from pre-trained models.
@@
Despite these advancements, existing methods still face significant challenges in providing reliable general-purpose protein understanding for scientific research applications. Firstly, although some studies acknowledge the critical role of a protein’s 3D structure in determining its properties and functions, proteins with directly available 3D structural information are rare. As a result, models must learn the relationship between 1D sequences and functional information while relying on limited 3D data, making it difficult to generate accurate property and function predictions. Secondly, existing protein-related instruction datasets overlook 3D structural information and provide limited coverage of protein properties and functions. This gap hinders the comprehensive evaluation of model reliability and generalizability in general-purpose protein understanding tasks. Lastly, the vast diversity of protein properties and functions presents a major challenge for accurate prediction. Using a single general-purpose model to predict a wide range of complex properties and functions is more challenging than fine-tuning multiple specialized models for specific tasks.
@@
To address these challenges, we propose a novel instruction tuning framework called Structure-Enhanced Protein Instruction Tuning (SEPIT) for general-purpose protein understanding. In this framework, we first design a structure-aware module within pLMs to obtain a protein sequence/structure-fused encoder, enabling support for different types of protein inputs (1D or 1D&3D). Following this, we implement a three-stage training pipeline. In stage 0 (warm-up stage), we train the protein sequence/structure-fused encoder using protein-text contrastive learning and structure denoising, allowing it to leverage a limited amount of structural information to enhance the understanding of large-scale sequence-only proteins. Then, we perform protein instruction tuning to equip LLMs for general-purpose protein understanding capability: in stage 1, we instill fundamental understanding of proteins into the model through protein caption instructions; in stage 2, we initialize mixture of experts (MoEs) through upcycling (Komatsuzaki et al., 2023), which allows the model to learn more complex and diverse protein properties and functions while building upon the foundational knowledge from stage 1, all without increasing the number of activated parameters. To comprehensively evaluate the reliability and generalizability of our proposed framework, we construct the largest protein instruction dataset to date, which covers the most diverse range of protein properties and functions, based on large-scale protein knowledge bases (Bairoch and Apweiler, 1997; Varadi et al., 2022). In summary, our key contributions are as follows:
  • We designed and integrated a structure-aware module into protein language models, enabling the models to process various types of protein inputs. This significantly improved embedding quality compared to vanilla sequence-only pLMs.
  • We constructed the largest and most extensive protein instruction dataset to date, addressing the gap in datasets for general-purpose protein understanding.
  • We developed a novel protein instruction tuning strategy that allows a single model to learn a broad and complex range of protein properties and functions. This was achieved by leveraging MoEs built upon foundational knowledge.
  • Using the proposed SEPIT framework and our curated protein instruction dataset, we demonstrated the feasibility of equipping large language models with general-purpose protein understanding capability.
2. Related Work
@@
In this section, we present the key related work pertinent to our study, focusing primarily on protein language models. Additional related topics including multimodal instruction tuning and learning with 3D structural information are discussed in Appendix A.
@@
Using context‐aware language models (Rosenfeld, 2000), protein sequences can be treated like sentences, with amino acids acting as individual words. Through pre-training on databases containing hundreds of millions of such protein sequences (e.g., UniRef (Suzek et al., 2015), BFD (Steinegger et al., 2019; Steinegger and Söding, 2018)), pLMs enable effective modeling and prediction of protein structures and functions (Hu et al., 2022). In earlier works (Alley et al., 2019; Strodthoff et al., 2020; Heinzinger et al., 2019), LSTM and its variants (Hochreiter and Schmidhuber, 1997; Yu et al., 2019; Huang et al., 2015; Krause et al., 2017) were utilized to model the dependencies between residues in single protein sequences. With the rise of the Transformers architecture (Vaswani et al., 2017), Transformers-based pLMs emerged. ESM-1b (Rives et al., 2021), leveraging the Transformers architecture along with a masking strategy for pretraining, significantly enhances the prediction accuracy for mutational effects, secondary structure, and long-range contacts. After this, ProtTrans (Elnaggar et al., 2022) released two auto-regressive models (Dai et al., 2019; Yang et al., 2019) and four auto-encoder models (Devlin et al., 2019; Lan et al., 2020; Clark et al., 2020; Raffel et al., 2020) pre-trained on protein sequence databases. Beyond merely focusing on single protein sequences, MSA Transformer (Rao et al., 2021) integrates multiple sequence alignments (MSA) of homologous proteins, providing a solid foundation for the success of AlphaFold2 (Jumper et al., 2021) and AlphaFold3 (Abramson et al., 2024). Moreover, ESM-2 (Lin et al., 2023) further scaled up pLMs, achieving protein structure prediction performance comparable to previous works without utilizing MSA information, and significantly reduced inference overhead (Lin et al., 2022). Additionally, there were other studies that attempted to incorporate additional knowledge into the pre-training of protein sequences. For example, ProteinBERT (Brandes et al., 2022) and OntoProtein (Zhang et al., 2022) incorporate Gene Ontology (GO) annotations directly into their protein embeddings. More recent methods—such as SaProt (Su et al., 2024) and ESM-3 (Hayes et al., 2025)—go further by fusing both sequence- and structure-based representations, thereby deepening the models’ grasp of protein characteristics. Although these protein language models produce rich, high-quality embeddings, they remain incapable of generating natural-language predictions about protein properties and functions.
3. Construction of Protein Instruction Dataset
@@
To endow LLMs with general-purpose protein understanding capabilities and evaluate their reliability and generalizability, in this paper, we construct a protein instruction dataset contains open-ended generation and closed-set answer tasks. For the open-ended generation subset, we mainly constructed it based on Swiss-Prot (Bairoch and Apweiler, 1997). We include almost all protein properties and functions contained therein (Function, Similarity, Subcellular location, Induction, Molecular Function, Biological Process, Cellular Component, Developmental Stage, Short Sequence Motif, Tissue Specificity, Activity Regulation, Pathway), and utilized ChatGPT (OpenAI et al., 2024) to aid in designing question templates based on the structured annotations. For the closed-set answer subset, we constructed it mainly based on the RCSB PDB (Berman et al., 2000). We follow the data organized by previous researchers (Guo et al., 2023) and select parts of their proposed Q&A samples that are highly related to protein properties and functions, filtering out other samples related to metadata (e.g. discovery time and discovery methods). We have also sampled some examples related to Enzyme Commission (EC) and Gene Ontology (GO) predictions (Gligorijević et al., 2021) for inclusion in the closed-set answer subset. More detailed information about the dataset, including statistical information and examples, is shown in Appendix B.
@@
Compared to previously proposed protein-text datasets (Xu et al., 2023; Fang et al., 2024; Wang et al., 2024c; Guo et al., 2023), our dataset offers several key advantages: First, our dataset contains the most comprehensive set of instructions, covering nearly all critical protein properties and function types found in databases (Bairoch and Apweiler, 1997). Second, with over 10 million instructions, along with an additional 5 million supplementary instructions from TrEMBL, our dataset is the largest of its kind. Third, by incorporating structural data, our dataset enhances prediction reliability and provides a more robust foundation for protein understanding. In summary, our dataset serves as a valuable resource for advancing research in general-purpose protein understanding.
Figure 1. (a) The model architecture of the SEPIT framework includes sequence/structure fused protein encoder, linear projector, and LLMs with MoEs modules, (b) example of instruction format.
4. Structure-Enhanced Instruction Tuning